### v2

@@ IV Results and Discussion
To optimize the Lable Propagation algorithm to achieve the best performance on automatic labeling, different proportions of the dataset were tried for annotation. The aim is to annotate as few pulses as possible. 2.5%, 5%, 7.5%, and 10% of the dataset were annotated, given that the entire database contains 1571 signals, and the proportion that gave the best results was evaluated. For each proportion of the dataset the precision, recall and F1 values were analyzed to decide. The results for the class 1 (class ”pulse with artifacts”) of these metrics are presented in the table II. Because the data are imbalanced, the results for the class 0 (class ”pulse without artifacts”) remain consistently good and don’t change much with different parameters. The best results are obtained for a proportion of 5% of the dataset. Indeed, as the proportion of annotated data increases, the distribution of classes becomes even more disparate. For 2.5% there are 17.3% of pulses with artifacts, and for 5% the proportion of pulses with artifacts is 18.1%. As the size of the annotated dataset increases, for 7.5% there are 16.4% of pulses with artifacts. For 10% of the dataset, 17.7% of pulses contain artifacts. All these values are summarized in the table I. If the classes are more unbalanced, this may have an influence on the algorithm, which will have greater difficulty in finding a constant pattern for propagating the labels. In the case of 10%, the proportion of pulses with artifacts is high, but the number of annotated pulses increases, and this may induce new data that is less representative of the overall data distribution, leading to poor generalization on unseen data.
TABLE I: Summary table containing the labeling portion and the imbalance rate
@@
TABLE II: Results for the class ”with artifacts” for different proportions of the dataset
@@
TABLE III: Results for the class ”with artifacts” for different sampling methods
@@
Figure 4: Confusion matrix and ROC curve for the Label Propagation algorithm with a KNN kernel with 7 neighbors, an oversampling method SMOTE and 5% of the dataset already labeled.
The various resampling methods presented in section II-C3 were applied. The results are shown in table III. The difference in results between undersampling and oversampling can be explained by the fact that undersampling will reduce the number of majority, this leads to loss of data and loss of information from this data. On the contrary, oversampling increases the number of values in the minority class, providing more data. In our case, SMOTE is the best oversampling method. SMOTE selects a minority class instance and identifies its k-nearest neighbors in the feature space. It then creates new synthetic examples along the line segments connecting the selected instance and its neighbors. By introducing these synthetic examples, SMOTE effectively increases the size of the minority class, making it comparable to the majority class and improving the performance of classifiers in handling imbalanced datasets.
TABLE IV: A comparison of the performance of different classifiers for the class ”with artifact”
Model Precision Recall F1 AUROC
KNN 0.89 0.95 0.92 0.98
MLP 0.76 0.97 0.85 0.99
Transformer 0.85 0.86 0.85 0.97
ConvolutionNN Baseline Classifier 0.86 0.83 0.84 0.95
@@
Figure 6: Accuracy curve during training and validation for the FCN model.
@@
Given the correct sampling method and the appropriate proportion of the dataset to be selected, the results of the Label Propagation algorithm were evaluated. The confusion matrix is shown on the left and the ROC curve on the right, on the Fig. 4. The ROC curve is plotted for each decision threshold. In the case of the label propagation algorithm, this represents the probability assigned to each instance for each class. For a 5% dataset, the number of pulses for the validation part is 1252 pulses. 1034 belong to the ”without artifacts” category and 218 to the ”with artifacts” category. The number of true positives and true negatives is higher than the number of false positives or false negatives. This indicates that the algorithm is correctly able to understand the model and apply it to unlabeled signals. However, the number of pulses detected as clean but actually containing artifacts (false negatives) is higher than the reverse (false positives). This is because the Label Propagation algorithm uses neighborhood information to propagate labels through the data network. This means that labels for samples close in feature space tend to be similar. However, in the case of pulses containing artifacts, these artifacts may be similar to certain features of other clean pulses, leading to incorrect label propagation. As a result, pulses containing artifacts may be incorrectly labeled as clean by the Label Propagation algorithm, leading to a higher number of falsely classified pulses. Having misclassified pulses is always a problem that is important in the medical field. This can lead to false alarms, if the pulse is not a clean pulse, or to misdetections. False alarms force hospital staff to make emergency visits due to outliers. These situations are exhausting, and not necessary as an additional burden on caregivers. The area under the ROC curve (AUROC) is 0.97. The closer the AUROC is to 1, the better the model’s performance. A high AUROC indicates that our model is able to discriminate well between positive and negative classes.
@@
After evaluating the Label Propagation algorithm, the performance of the different types of classifiers, presented in section II-D, were assessed. First, dealing with the imbalanced classes by oversampling using Adaptive Synthetic (ADASYN) algorithm. Cross-validation was employed to ensure accurate model prediction and assess the reliability of the machine learning algorithms. The results of the 5-fold cross-validation on different classifiers are presented in Fig. 5, which shows a precision comparison using a box plot. Each blue dot represents the performance of an individual fold in the cross-validation. The figure indicates that KNN and SVC (with a kernel of ’rbf’) are the top-performing classifiers, with median precision rates above 90%. However, KNN shows a slightly better and more consistent performance than SVC. This observation is highlighted by the broader range of variability for SVC, as indicated by the whiskers on the box plot, compared to KNN. To sum up, KNN and SVC are the top classifiers, but KNN is the more reliable and stable solution. KNN is also the best classifier compared to classifiers that use neural networks such as MLP and Transformers. Table IV shows the different performances of the neural networks classifiers: MLP classifier, Transformer, ConvolutionNN VS KNN classifier. MLP consists of 3 hidden layers with 500 neurons for each hidden layer. Its macro average accuracy (calculates the accuracy for each class individually and then computes the average accuracy across all classes) is 0.88, compared with 0.94 for KNN. In our case, using a complex model like MLP could lead to overfitting, as the model may have a high capacity relative to the amount of data available. In addition, training an MLP can be computationally expensive, especially with larger architectures and limited computational resources. Transformers, especially large ones like BERT (Bidirectional Encoder Representations from Transformers), have a high computational complexity and require significant computational resources for training and inference. Like the MLP classifier, Transformers works best on larger datasets because it needs a lot of data for the training part, otherwise the model has a greater capacity than the limited number of data and the risk is overfitting. Generally speaking, in the medical field, Transformers excel in natural language processing tasks [34]. They can learn complex relationships and patterns within the text, making them suitable for tasks involving medical text classification and understanding.
@@
For the last classifier, experiments were conducted with a Fully Convolutional Network (FCN). FCN is a neural network architecture designed for semantic segmentation, producing dense pixel-wise predictions. It consists of convolutional layers without fully connected layers, enabling it to handle images of any size and preserve spatial information. Using FCN for time series classification involves adapting the fully convolutional architecture to process one-dimensional time series data. Instead of working with two-dimensional images, the FCN is applied to sequences of data points. The temporal convolutional layers capture temporal patterns and dependencies in the time series, and the decoding path with transposed convolutional layers helps to produce dense predictions for each data point in the sequence, enabling accurate time series classification [35]. One key benefit is that FCN eliminate the need for manual feature engineering, as they can directly learn relevant features from raw time series data. This streamlines the classification process and saves time and effort in designing handcrafted features. Additionally, FCN enable end-to-end learning, optimizing feature representations and classification jointly, which can lead to improved performance. The flexibility of FCN with input size allows them to handle time series data of varying lengths without requiring resizing or padding, making them suitable for irregularly sized data. Moreover, FCN produce dense predictions for each time step, capturing fine-grained temporal patterns and enhancing the informativeness of classification results. Experimentally, during FCN training, it is evident that the process takes longer than other approaches. However, its performance is not comparable to those methods, mainly due to its lower accuracy. Moreover, the Fig. 6 shows that both the validation and training accuracy display fluctuations, indicating a lack of stability in its performance.
V Conclusion
In conclusion, this study delved into utilizing semi-supervised label propagation (LP) methods for artifact classification within PPG signals, especially in scenarios characterized by imbalanced class distributions. To validate the effectiveness of this approach, we compared the results from the semi-supervised LP method against those obtained using traditional supervised learning algorithms, MLP, and Transformer-based models. Based on the experimental results, with oversampling method, the Label Propagation classifier performs better than neural network classifiers, MLP, FCN, and Transformers, achieving precision, recall, and F1 scores of 91%, 90%, and 90%, respectively. In particular, the performance of Label Propagation is comparable to that of KNN (k=3), which attains a higher recall score of 95% and an F1 score of 92%. However, the label propagation performs more consistently with a balanced precision-recall trade-off, whereas the KNN classifier struggles in that aspect. For improvements and future work, some directions can be given:
@@ V Conclusion
  • Improving the preprocessing: In section II-C2, the problem of segmentation has already been mentioned. Some steps can be added to the preprocessing part. First, adaptive filtering techniques can be used to attenuate artifacts without affecting the signal. Signal quality can be improved through noise reduction methods, such as singular value decomposition (SVD). To enhance the efficiency of the statistical analysis algorithm, alternative segmentation approaches can be employed. Peak or minima detection can be improved by employing derivative-based algorithms. A CNN model can also be used to detect peaks, known for its pattern recognition capabilities.
  • Machine learning approaches: Adapting different methods for tackling labeling limitation like self-supervised or unsupervised learning approaches.
  • Data augmentation: Exploring data augmentation methods to increase data availability for improving the classifier’s generalization.
Overall, the use of label propagation techniques for labeling and classifying artifacts detection in PPG signals presents a promising direction for more accurate and reliable healthcare monitoring systems.
@@ Acknowledgment
The clinical data used in this study were generously provided by the Research Center of the CHU Sainte-Justine Hospital, University of Montreal. The authors thank Dr. Sally Al Omar, Dr. Michael Sauthier, Edem Tiassou for their data support of this research and Kévin Albert for annotating the data.

### v3

@@ V Results and Discussion
Different proportions of the dataset were tried for annotation to optimize the Lable Propagation algorithm and achieve the best performance on automatic labeling. The aim is to annotate as few pulses as possible. 2.5%, 5%, 7.5%, and 10% of the dataset were annotated, given that the entire database contains 1571 signals, and the proportion that gave the best results was evaluated. For each proportion of the dataset, the precision, recall, and F1 values were analyzed to decide. The results for class 1 (class ”pulse with artifacts”) of these metrics are presented in the table IV. Because the data are imbalanced, the results for class 0 (class ”pulse without artifacts”) remain consistently good and don’t change much with different parameters. The best results are obtained for a proportion of 5% of the dataset. Indeed, as the proportion of annotated data increases, the distribution of classes becomes even more disparate. For 2.5% there are 17.3% of pulses with artifacts, and for 5%, the proportion of pulses with artifacts is 18.1%. As the size of the annotated dataset increases, for 7.5% there are 16.4% of pulses with artifacts. For 10% of the dataset, 17.7% of pulses contain artifacts. All these values are summarized in the table III. If the classes are more unbalanced, this may influence the algorithm, which will have greater difficulty in finding a constant pattern for propagating the labels. In the case of 10%, the proportion of pulses with artifacts is high, but the number of annotated pulses increases, and this may induce new data that is less representative of the overall data distribution, leading to poor generalization on unseen data.
TABLE III: Summary table containing the labeling portion and the imbalance rate
@@
TABLE IV: Results for the class ”with artifacts” for different proportions of the dataset
@@
TABLE V: Results for the class ”with artifacts” for different sampling methods
@@
Figure 5: Confusion matrix and ROC curve for the LP algorithm with a KNN kernel with 7 neighbors, an oversampling method SMOTE, and 5% of the dataset already labeled.
The various resampling methods presented in section III-C3 were applied. The results are shown in table V. First, we can see that despite the class imbalance, the algorithm manages to detect the artifacts for the training part and that the scores are correct (96% precision, 82% recall, and 89% F1). However, when we apply a resampling method, the results between the scores are more balanced. This results in a more robust algorithm. The difference in results between undersampling and oversampling can be explained by the fact that undersampling will reduce the number of majority, which leads to loss of data and loss of information from this data. On the contrary, oversampling increases the number of values in the minority class, providing more data. In our case, SMOTE is the best oversampling method. SMOTE selects a minority class instance and identifies its k-nearest neighbors in the feature space. It then creates new synthetic examples along the line segments connecting the selected instance and its neighbors. By introducing these synthetic examples, SMOTE effectively increases the size of the minority class, making it comparable to the majority class and improving the performance of classifiers in handling imbalanced datasets. However, it is important to consider that resampling can potentially cause issues such as overfitting. It is important to monitor the model’s performance after oversampling to detect any signs of overfitting or other potential issues. Our model showed no signs of overfitting, and resampling was very important for training a well-balanced classifier in the case of the imbalanced dataset.
TABLE VI: A comparison of the performance of different classifiers for the ”with artifact” class. The oversampling method chosen is SMOTE for the LP model and ADASYN for the other models. In addition, 5% of the dataset is already annotated.
Model Precision Recall F1 MCC Kappa CSI AUROC
LP 0.91 0.90 0.90 0.89 0.89 0.82 0.98
KNN 0.89 0.95 0.92 0.90 0.90 0.85 0.98
MLP 0.76 0.97 0.85 0.83 0.82 0.74 0.99
Transformer 0.85 0.86 0.85 0.82 0.82 0.74 0.97
FCN 0.86 0.83 0.84 0.82 0.82 0.74 0.95
@@
Figure 7: The comparison of CNN performance during training (Train) and validation (Val) for loss, AUROC, precision, and recall.
@@
Given the correct sampling method and the appropriate proportion of the dataset to be selected, the results of the LP algorithm were evaluated. The confusion matrix is shown on the left and the ROC curve on the right, on the Fig. 5. The ROC curve is plotted for each decision threshold. In the case of the LP algorithm, this represents the probability assigned to each instance for each class. For a 5% dataset, the number of pulses for the validation part is 1252. 1034 belong to the ”without artifacts” category, and 218 belong to the ”with artifacts” category. The number of true positives and true negatives is higher than the number of false positives or false negatives. This indicates that the algorithm can understand and apply the model to unlabeled signals. However, the number of pulses detected as clean but containing artifacts (false negatives) is higher than the reverse (false positives). The LP algorithm uses neighborhood information to propagate labels through the data network. This means that labels for samples close in feature space tend to be similar. However, in the case of pulses containing artifacts, these artifacts may be similar to certain features of other clean pulses, leading to incorrect label propagation. As a result, pulses containing artifacts may be incorrectly labeled as clean by the LP algorithm, leading to a higher number of falsely classified pulses. Misclassified pulses are always an important problem in the medical field. This can lead to false alarms if the pulse is not a clean pulse or to misdetections. False alarms force hospital staff to make emergency visits due to outliers. These situations are exhausting and not necessary as an additional burden on caregivers. Presented in Table VI, the three additional scores, MCC, Kappa, and CSI, support this explanation. Indeed, these three values must be close to 1 to indicate a good algorithm prediction. The closer the score is to 0, the more random the algorithm’s prediction. We note that the MCC, CSI, and Kappa values demonstrate a strong prediction of the LP algorithm. These scores allow us to validate the algorithm’s correct performance. The AUROC is 0.98. The closer the AUROC is to 1, the better the model’s performance. A high AUROC indicates that our model can distinguish between positive and negative classes.
@@
After evaluating the LP algorithm, the performance of the different types of classifiers, presented in section III-D, were assessed. First, dealing with the imbalanced classes by oversampling using the ADASYN algorithm. Cross-validation was employed to ensure accurate model prediction and assess the reliability of the machine-learning algorithms. The results of the 5-fold cross-validation on different classifiers are presented in Fig. 6 shows a precision comparison using a box plot. Each blue dot represents the performance of an individual fold in the cross-validation. The figure indicates that KNN and SVC (with a kernel of ’rbf’) are the top-performing classifiers, with median precision rates above 90%. However, KNN shows a slightly better and more consistent performance than SVC. This observation is highlighted by the broader range of variability for SVC, as indicated by the whiskers on the box plot, compared to KNN. To sum up, KNN and SVC are the top classifiers, but KNN is the more reliable and stable solution. KNN is also the best classifier compared to classifiers that use neural networks such as MLP and Transformers. Table VI shows the different performances of the neural networks classifiers: MLP classifier, Transformer, Fully Convolutional Network (FCN) VS KNN classifier. MLP consists of 3 hidden layers with 500 neurons for each hidden layer. Its macro average accuracy (calculates the accuracy for each class individually and then computes the average accuracy across all classes) is 0.88, compared with 0.94 for KNN. In our case, using a complex model like MLP could lead to overfitting, as the model may have a high capacity relative to the amount of data available. In addition, training an MLP can be computationally expensive, especially with larger architectures and limited computational resources. Transformers, especially large ones like BERT (Bidirectional Encoder Representations from Transformers), have a high computational complexity and require significant computational resources for training and inference. Like the MLP classifier, Transformers works best on larger datasets because it needs a lot of data for the training part. Otherwise, the model has a greater capacity than the limited data, and the risk is overfitting. Generally speaking, in the medical field, Transformers excel in natural language processing tasks [37]. They can learn complex relationships and patterns within the text, making them suitable for medical text classification and understanding tasks.
Additionally, The LP algorithm demonstrates consistent and balanced performance in precision and recall, as evidenced by the experiment results in Fig. 5 and Table VI. The confusion matrix in Figure 5 shows that the LP algorithm with a KNN kernel (7 neighbors) achieves high accuracy, correctly classifying the majority of positive and negative cases, leading to a precision of 0.91 and a recall of 0.90. The ROC curve with an area of 0.98 further indicates the model’s ability to distinguish between classes. Table 6 reinforces these findings by comparing the LP algorithm to other classifiers, where LP maintains competitive precision and recall values while achieving high F1 (0.90) and MCC (0.89) scores. These results highlight that the LP algorithm, combined with SMOTE for oversampling and leveraging 5% of already labeled data, effectively balances precision and recall, ensuring robust performance in classifying the ”with artifact” class.
@@
For the last classifier, experiments were conducted with an FCN model. FCN is a neural network architecture for semantic segmentation, producing dense pixel-wise predictions. It consists of convolutional layers without fully connected layers, enabling it to handle images of any size and preserve spatial information. Using FCN for time series classification involves adapting the fully convolutional architecture to process one-dimensional time series data. Instead of working with two-dimensional images, the FCN is applied to sequences of data points. The temporal convolutional layers capture temporal patterns and dependencies in the time series, and the decoding path with transposed convolutional layers helps to produce dense predictions for each data point in the sequence, enabling accurate time series classification [38]. One key benefit is that FCN eliminates manual feature engineering, as they can directly learn relevant features from raw time series data. This streamlines the classification process and saves time and effort in designing handcrafted features. Additionally, FCN enables end-to-end learning, optimizing feature representations and classification jointly, which can lead to improved performance. The flexibility of FCN with input size allows them to handle time series data of varying lengths without requiring resizing or padding, making them suitable for irregularly sized data. Moreover, FCN produces dense predictions for each time step, capturing fine-grained temporal patterns and enhancing the informativeness of classification results. Experimentally, during FCN training, it is evident that the process takes longer than other approaches. However, its performance is not comparable to those methods, mainly due to its lower accuracy.
For training MLP, FCN, and Transformer, we use the binary cross-entropy loss as follows:
$L_{BCE}=-\frac{1}{n}\sum_{i=1}^{n}(Y_{i}\times log(\hat{Y}_{i})+(1-Y_{i})\times log(1-\hat{Y}_{i}))$ (13)
We use the Adam optimizer and early stopping to deal with the overfitting. We use GridSearchCV to fine-tune the hyper-parameters, balancing the best combination and computation time. Only certain hyper-parameters typically affect a neural network’s accuracy, specifically the number of hidden layers, nodes in each hidden layer, and the learning rate [39]. By focusing on these, grid search effectively optimizes all parameters simultaneously, allowing for quick model training. Grid search also offers straightforward parallelization and flexible resource allocation, which other approaches lack [40].
Fig. 7 provides a comprehensive view of the model’s performance over time. The improvements in metrics like loss, AUC, precision, and recall suggest that the model is learning and improving its performance with each epoch. The consistent trends between training and validation data indicate that the model is generalizing well and not overfitting significantly. However, we can see the fluctuation between precision and recall; it can be confirmed that FCN can not deal with the imbalanced classes from the nature of the data.
So, compared with previous studies, the artifact classification algorithm we have implemented has the advantage of having a faster execution time on large volumes of data compared to EMD or wavelet denoising, for example. It exploits the intrinsic relationships between the data rather than decomposing each signal individually. It also has the advantage of being easily generalizable to other signals since no additional parameters are required.
VI Limitations and future work
@@
This study delved into utilizing semi-supervised LP methods for artifact classification within PPG signals, especially in scenarios characterized by imbalanced class distributions. The study showed us that our model is sensitive to data volume, and its improvement is limited as data volume increases. One future objective is to improve our model, particularly in feature detection. To augment the capability of our model, we can add some steps in the preprocessing part. In section III-C2, the segmentation problem has already been mentioned. First, adaptive filtering techniques can attenuate artifacts without affecting the signal. Signal quality can be improved through noise reduction methods, such as singular value decomposition (SVD). Alternative segmentation approaches can be employed to enhance the efficiency of the statistical analysis algorithm. Peak or minimum detection can be improved by employing derivative-based algorithms. A CNN model can also detect peaks, known for its pattern recognition capabilities [41]. Implementing alternative segmentation approaches and employing a CNN model for peak detection in PPG signals may face challenges in parameter tuning, validation, and network architecture design. However, careful optimization and validation of these approaches can improve the algorithm’s accuracy and reliability.
Exploring data augmentation can also be a method to tackle the model’s sensitivity to data volume. The authors in [42] found that using data augmentation allowed them to better handle the unbalanced class problem for binary or multiclass classification. These authors also draw a parallel with ensemble learning methods, which are new hybrid methods that are more robust against unbalanced data. In [43, 44], the authors put into practice the use of RUSBoost for epilepsy seizure detection and schizotypy classification. Compared to classical models, the ensemble model performed very well, making it a candidate for feature classification. Investigating ensemble learning models may help to better handle class imbalances and classification tasks. Exploring self-supervised or unsupervised learning methods could be considered for future work to address the labeling challenge we encountered without relying on manual annotations.
VII Conclusion
This study explored applying semi-supervised LP methods for artifact classification in PPG signals, addressing challenges posed by imbalanced class distributions. Comparative analysis with traditional supervised learning algorithms, MLP, and Transformer-based models demonstrated the superior performance of the LP classifier. This algorithm can enhance the overall quality of PPG signals in artifact classification by dynamically adapting to the specific characteristics of the dataset. It becomes more adaptable to variations in PPG signals caused by different types of motion artifacts. The improved balance between precision and recall indicates more robust classifier performance, which is critical for real-life medical applications.
Overall, this model holds promise for enhancing healthcare monitoring systems, with potential applications in ECG and arterial blood pressure signal analysis.
@@ Acknowledgment
The clinical data used in this study were generously provided by the Research Center of the CHU Sainte-Justine Hospital, University of Montreal. The authors thank Dr. Sally Al Omar, Dr. Michael Sauthier, and Edem Tiassou for their data support of this research and Kévin Albert for annotating the data.